Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2025 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 5     Entry time: 30 Jan 2025, 23:08
Author: Chan Kim 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 11.0 keV, up to 178 pulses at 564 kHz, 650 uJ.

Optical laser delivery

  • illumination and pp lasers

Achievements/Observations

  • Performed detector and ice detection program.

Issues

  • Run Controller was crashing several times during the day. DOC was notified. They discovered that the run controller server was down and fixed the problem by restarting it.
  • PP laser was not available during the day until ~ 18:30
  • Some troubles with and homing of SmarAct stages after update from to the current control version. DOC was requested to roll back to the latest stable release.
ELOG V3.1.4-7c3fd00